General Meeting Summary 5/12/22

From Cuawiki
Jump to: navigation, search

PARTICIPANTS: Vladimir Berdnikov, Charles Hyde, Aaron Brown, Abishek Karki, Arshak Asaturyan, Bogdan Wojtsekhowski, Brad Sawatzky, Carlos Munoz-Camacho, Deb, Erik Wrightson, Hamlet Mkrtchyan, Josh Crafts, Julie Roche, Mark Jones, Marie Boer, Paulo Medeiros, Pramita Tiwari, Rolf Ent, Steven Lassiter, Vardan Tadevosyan, Tanja Horn, Petr Stepanov, Josh Crafts

CRYSTAL WRAPPING (Vladimir, Petr, Josh, Hamlet, Vardan, Abishek, Zheng Huang)

  • 7 members (CUA, AANL, MSU, UIC)
  • working in multiple teams of 1-2 people throughout the day
  • aim to wrap 1100 PWO for NPS stacking by end of May
  • present documentation:
  • handwritten sheets
  • crystal ID - cleaned - initials - date
  • online spreadsheet (electronic documentation) in preparation
  • crystals stacked on shelves - estimate to need two more shelves
  • Wrapping details
  • good wrapping is the one with a small gap (20% have ~mm gap and 80% have ~um gap)
  • overlap of tedlar should be avoided as it impacts the stacking
  • avoid real crinkling of tape/tedlar as it would complicate stacking - currently don't have any real crinkling
  • Two additional shelves are now on site - have to be assembled
  • Action items:
  • avoid any overlap of tedlar - if there is one cut the tedlar to make sure that there is a small gap
  • make sure that keep 2cm space on both ends of the crystal
  • assemble the two additional shelves


  • Beam schedule will come out soon
  • NPS run group 1 experiments will run 2023/24

MECHANICAL (Paulo, Steven L.)

  • finished platform
  • currently working on rail to base of detector
  • pre-installation nearly completed - some items not done
  • next week take apart again for experiments starting in June
  • what was not done: reinforcement on pivot (has to do at beginning of next SAD), has to be included in schedule
  • note that schedule is tight next year and will need much technical support to make the schedule
  • Action item:
  • send updated schedule


  • Controls and Monitoring
  • Ansys thermal analysis
  • central crystals at 22degC, peripheral at 10degC
  • studies show that the main constraint is the ambient temperature (20deg assumed here) --> it follows that one cannot achieve constant temperature that is lower than that with the NPS adopted cooling system, i.e., cooling from the outside
  • note: NPS is assuming ambient temperature, so all fine - mainly need to focus on maintaining ambient temperature in hall and measure the temperature
  • another note: PANDA is cooling to -25degC, but has a different cooling system where a cooling fluid is circulated between the individual crystals
  • High voltage cables
  • gluing the cables to the connectors - no difference in performance observed

Action items:

  • continue NPS activities (thermometry, readout, etc.)
  • prepare publication of thermal analysis

DAQ/SOFTWARE (David H. per email)

  • The main updates from last week's short DAQ/software meeting were from Brad regarding the status of slow controls (everything in good shape except the LED control GUI, which still has some work to do -- but this is not a problem as we can control from the terminal for now) and getting things ready for the move to EEL108.
  • Julie and Pramita are working on online displays based on the pionLT SHMS calorimeter mode 10 data as are Oliver and myself.
  • Action item:
  • Present a comprehensive update at the next full NPS meeting


  • Coincidence Parity (Bogdan)
  • number of PWO crystals
  • radiation estimation at location of crystals
  • Neutron DVCS (Carlos)
  • unique measurement of cross sections
  • can also measure pi0 - maybe emphasize more?
  • Action items:
  • circulate proposal to NPS
  • comment period until Saturday 5/14 end of day - if any objection to list "and the NPS collaboration" on the proposal indicate that clearly
  • if no objections are received both proposals can list "and the NPS collaboration" in the submission